If f'(x) = 9 x+ 12 and f(5) = -11, find f(x).
If f'(x) = 9 x+ 12
I have f(x) = (9/2)x^2 + 12x + C
But I am not sure where to go from there
I am completely lost..
THANK YOU

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

If f ′(x) = 9 x+ 12 and f(5) = −11, find f(x).
If f ′(x) = 9 x+ 12
I have f (x) = 9x^2/2 + 12x + C
----
Solve for "C" using f(5) = -11
----
-11 = (9/2)(5)^2+12(5) + C
-71 = (9*25/2 +C
-71 = 112.5
C = -183.5
---
f(x) = (9/2)^2 + 12x - 183.5
-------------
Cheers,
Stan H.
-------------

Answer by solver91311(24713)



I did a little editing on your post for readability's sake but did not change the meaning in any way.

You certainly aren't completely lost since everything you posted was correct. I'm only going to give you a hint at this point. Since you integrated the derivative, the result was the family of functions that have that derivative. So what does the fact that tell you when you know that ?

John

My calculator said it, I believe it, that settles it